OpenStudy (anonymous):

Donna opens a certificate of deposit (CD) with $2,000. The bank offers a 3% interest rate. If the account compounds quarterly, which of the following equations represents the future value of the account, after 1 year? A. A = 2000 (0.03/ 4) ^4

OpenStudy (anonymous):

B. A = 2000 ( 0.03/4) C. A = 2000 (1+ 0.03)^4 D. A = 2000 (1 + 0.03/4)^4

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

I'm sorry, I'm not good with financial mathematics